OpenStudy (anonymous):

Solve. Y=2x+11 Y=x=8

OpenStudy (p0sitr0n):

y=x=8

OpenStudy (anonymous):

OHHHH I MEANT FOR THE SECOND ONE Y=X+8

OpenStudy (p0sitr0n):

compare y=y 2x+11 = x+8 x = -3 y= 2(-3) +11 y = 5